40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 28 21:51:52 2007, tobin, Configuration, PSL, ISS 
    Reply  Thu Nov 29 15:17:53 2007, rob, Configuration, PSL, ISS 
       Reply  Mon Dec 3 22:20:09 2007, tobin, Configuration, PSL, ISS 
Message ID: 137     Entry time: Wed Nov 28 21:51:52 2007     Reply to this: 141
Author: tobin 
Type: Configuration 
Category: PSL 
Subject: ISS 
I replaced the front-end differential receivers for the ISS's "inner-loop" sensor and monitor diode inputs with lower-noise THS4131's (formerly THS4151's). I verified operation by taking the transfer function from the "PD+" and "PD-" inputs (separately) to the testpoint following the differential receiver; the surgery appears successful.

I measured the dark spectra at the ISS's DC PD BNC ports and found a noise floor of ~ 16 nV/rtHz, compared with a floor of ~ 22 nV/rtHz last week. This seems to add up, assuming the DC PD port has 0dB gain: the 4131 has a rated noise of 1.3 nV/rtHz and the 4151 a noise floor of 7.6 nV/rtHz, a difference of 6 nV/rtHz. The other change made in that time was to add a larger power supply bypass capacitor in the PD.

There are two of the old 4151 chips still on the ISS board on the two "outer-loop" channels that we don't use. If I dig up any more 5131's I will replace these too for completeness.

There is currently no light on the ISS diodes; I'm not sure where it's intended to come from.
ELOG V3.1.3-